(2/(x+4)) - (3/(x-1))>=0
Is the answer (-4,1)?
Thank you,
Alexus

Answer by Prithwis(166)

(2/(x+4)) - (3/(x-1))>=0
=> [2(x-1) - 3(x+4) ]/(x+4)(x-1) >= 0
=> [-x-14]/(x+4)(x-1) >= 0 .........(1)
It seems you are looking for intervals.
The Boundary points are x=-4, x=1, x = 14
For x< -14, (1) is positive
For x in (-14, -4), (1) is negative
For x in (-4,1), (1) is positive
For x > 1, (1) is negative
The answer would be x belongs to (-infinity, -14] U (-4, 1)